# OBELiX: A Curated Dataset of Crystal Structures and Experimentally Measured Ionic Conductivities for Lithium Solid-State Electrolytes

### v1

@@ 4.1 Baselines
We chose two standard machine learning models, a random forest (RF) and a multilayer perceptron (MLP), to train on the full dataset.
We also tested five widely adopted graph neural networks developed specifically for materials science applications, PaiNN(Schütt et al., 2021), SchNet(Schütt et al., 2017), M3GNet(Chen & Ong, 2022), SO3Net(Schütt et al., 2023), and CGCNN(Xie & Grossman, 2018) on the subset of the dataset that contains CIF files. These graph-based models, where each node represents an atom, effectively capture atomic interactions while preserving molecular invariance, enabling accurate material property predictions when trained on large datasets (Liu et al., 2024).
  • PaiNN enforces E(3)-equivariance, enabling accurate modeling of atomic interactions and force predictions.
  • SchNet learns continuous filter representations, making it effective for capturing atomic environments.
  • M3GNet integrates message passing with three-body interactions, improving property predictions for crystalline materials.
  • SO3Net leverages spherical harmonics to enhance equivariant representations for molecular and solid-state systems.
  • CGCNN models crystal structures directly as graphs, making it a strong baseline for learning structure-property relationships.
@@ 4.2 Setup
To optimize the training process and assess the stability of the models, we implemented a 5-fold cross-validation strategy. For hyperparameter optimization, we employed a grid search strategy across a predefined space of 100 randomly sampled hyperparameter sets for each model. This number was selected to strike a balance between comprehensive exploration of the hyperparameter space and computational feasibility. In the case of RF and MLP where training is extremely fast; all hyperparmeter sets were tested. The hyperparameter space was carefully designed for each model based on its unique architecture and requirements (see Table 4 in the SI for a complete list). For example, PaiNN’s search space included parameters such as the cutoff distance, number of interactions, and batch size.
@@
We computed the mean absolute error (MAE) to evaluate the performance of each configuration. Specifically, the average validation MAE across all folds in the cross-validation process was used to assess each setup’s effectiveness. The hyperparameter set that achieved the lowest average validation MAE was selected as the best-performing configuration.
After choosing the best hyperparameters, each model was retrained on the entire training set and evaluated on the test set. This ensured that the results reflected the models’ optimal performance. A detailed table of the selected hyperparameters for each model is included in the SI (Table  4).
@@
Pretraining can enhance model performance by initializing weights with knowledge from a large dataset, which is then fine-tuned on a smaller, task-specific dataset. During fine-tuning, the pretrained model is adapted to the target dataset. We pretrained PaiNN and SchNet on the Materials Project with a band gap prediction task. In this case we fixed the trained representation (PaiNN or SchNet) and trained the output model (an MLP followed by a pooling layer) on OBELiX. For M3GNet and CGCNN we use pretrained models that were available on their public repositories. The M3GNet model was trained on formation energy per atom whereas CGCNN was trained on Fermi energy both from the Materials Project. As recommended in their respective documentation, we fine-tuned the models by training all model parameters starting from the trained models.
Table 2: Benchmarking of various ML models with and without pretraining.
@@
4.3 Discussion
Table 2 presents our benchmarking results. The fine-tuned version of each model is denoted with the prefix ”p-,” such as p-PaiNN for the fine-tuned PaiNN model.
The two simple models, RF and MLP outperform all 3D geometric models both in the cross-validation and the test performance. There are two factors that could explain this result. First, the RF and the MLP used the full training set of 478 structures while the other models were limited to the subset of 254 entries that have CIF files. The performance of the RF and the MLP on the subset of the training set that has CIF files is indicated in parentheses in Table 2 in order to be compared fairly with other models. Second, the geometric models use crystal information to infer properties of the crystal, but they not properly handle partial occupancies which, as discussed before, are very common in SSE materials and are present in about 3/4 of our CIF files. In order to use these models on our dataset we rounded occupancies to the nearest integers which can lead to important changes in the composition.
To partly verify the above claim that dataset size and the presence of partial occupancy can explain the increased performance of the simple models, we retrained them on the subset of entries that have CIF files only. Doing so, the MAE of the MLP dropped to 3.3 while that of the RF was maintained at 1.86. Data size does seem to have a significant impact on the MLP and may explain the difference in performance between that model and the larger models. Random forest still performs well even given less data. Rounding compositions to the nearest integer on the other hand, had little effect on both the RF and the MLP. Rounding compositions is similar to rounding site occupancy, but it does not have exactly the same effect. Nevertheless, it indicates that the absence of partial occupancy likely does not explain the difference in performance between the simple models and the more complex ones.
The 3D geometric models not only performed poorly compared to simple ML models using less structural information, but their performance on the test set was barely better or sometimes worse than predicting the median of the training set. This shows that these large models can easily overfit small experimental datasets which was also observed in other studies (Fung et al., 2021). The relatively large difference between the average validation MAE and the test set also shows the importance of carefully building leakage-free test sets.
Pretraining of 3D geometric models offers some marginal improvements for PaiNN, SchNet and CGCNN. As mentioned in Section 4.2, the pretraining of PaiNN and SchNet restricts the trainable model size which may reduce accuracy while increasing generalizability. This would explain their slightly higher validation MAE and lower training MAE. A better pretrained representation might compensate for the reduced expressivity and increase both accuracy and generalizability, but a much more in-depth analysis of the possible pretraining labels and datasets would be required. In the case of CGCNN and M3GNet which were fine-tuned by allowing all model parameters to change, it is possible that the pretraining property used for CGCNN was “closer” (or more relevant) to ionic conductivity which allowed it to stay in the same weight “basin” and take advantage of the pretrained model’s generalizability. It is important to bear in mind that the variability of the prediction accuracy is high in this small data regime as illustrated by the validation MAEs’ standard deviations and that much of the difference between these models falls within that variability. Therefore one must be careful when interpreting the results.
Ultimately, the poor performance of models in Table 2 highlights the need for domain specific models tailored for small data regimes.
5 Conclusion and Outlook
In this paper, we presented, OBELiX, a domain-expert-curated dataset of 599 materials with experimental room temperature ionic conductivities, including 321 structures with full crystallographic information. We gathered these materials from existing databases and manually extracted data from the literature to build a consistent, easy-to-access database of solid-state electrolyte materials. We benchmarked several ML models and found that (1) existing architectures likely over-fit and are unable to perform well on our carefully engineered test set, and (2) current geometric models cannot properly handle partial occupancies, which are very common in SSE materials.
We hope that OBELiX will inspire the development of new ML methods specific to this task and serve as a reference point to train and test ionic conductivity models for the ML and computational materials science community in general, ultimately advancing solid-state battery technology.
6 Data Availability
All data is freely available on our public repository222github.com/NRC-Mila/OBELiX as a single csv or xlsx file accompanied by a set of 321 CIF files, including 291 with added random noise.
Appendix A Supplemental Information
Table 3 presents the cross-validation and test MAEs for the models trained on randomized atomic positions. The models were trained on the randomized data and tested on the original CIFs. Table 4 shows the best hyperparameter sets for each 3D model presented in Table 2
Table 3: Performance of the 5 geometric models on the set that includes randomized atomic positions.
Model Cross-validation MAE Test MAE
PaiNN 2.03 $\pm$ 0.27 2.95
SchNet 1.99 $\pm$ 0.22 2.78
M3GNet 1.83 $\pm$ 0.29 2.91
SO3Net 1.98 $\pm$ 0.23 2.79
CGCNN 1.94 $\pm$ 0.42 2.95
Table 4: The selected hyperparameters for the models.
@@ Appendix A Supplemental Information
target_property ”graph”
readout_type ”set2set”

### v2

@@ 4.1 Baselines
To evaluate the performance of ML models on OBELiX, we tested five widely adopted graph neural networks developed specifically for materials science applications, PaiNN [25], SchNet [27], M3GNet [7], SO3Net [26], and CGCNN [32] on the subset of the dataset that contains CIF files. These graph-based models, where each node represents an atom, effectively capture atomic interactions while preserving molecular invariance, enabling accurate material property predictions when trained on large datasets [18]. On the full dataset, where atomic positions are not always available, we also tested two standard machine learning models, a random forest (RF) and a multilayer perceptron (MLP).
@@ 4.2 Setup
To optimize the training process and assess the stability of the models, we implemented a 5-fold cross-validation strategy. For hyperparameter optimization, we employed a grid search strategy across a predefined space of 100 randomly sampled hyperparameter sets for each model. This number was selected to strike a balance between comprehensive exploration of the hyperparameter space and computational feasibility. In the case of RF and MLP where training is extremely fast; all hyperparmeter sets were tested. The hyperparameter space was carefully designed for each model based on its unique architecture and requirements (see Table S2 in the SI for a complete list). For example, PaiNN’s search space included parameters such as the cutoff distance, number of interactions, and batch size.
@@
We computed the mean absolute error (MAE) between the predicted and the measured ionic conductivities to evaluate the performance of each configuration. Specifically, the average validation MAE across all folds in the cross-validation process was used to assess each setup’s effectiveness. The hyperparameter set that achieved the lowest average validation MAE was selected as the best-performing configuration. After choosing the best hyperparameters, each model was retrained on the entire training set and evaluated on the test set. A detailed table of the selected hyperparameters for each model is included in the SI (Table  S2).
@@
Pretraining can enhance model performance by initializing weights with knowledge from larger datasets and related tasks, which is then fine-tuned on a smaller, task-specific dataset. We pretrained PaiNN and SchNet on the Materials Project with a band gap prediction task. In this case we fixed the trained representation (PaiNN or SchNet) and trained the output model (an MLP followed by a pooling layer) on OBELiX. For M3GNet and CGCNN we use pretrained models that were available on their public repositories. The M3GNet model was trained on formation energy per atom whereas CGCNN was trained on Fermi energy both from the Materials Project. As recommended in their respective documentation, we fine-tuned the models by training all model parameters starting from the trained models.
Figure 4: Benchmarking of various ML models. The same data is tabulated in Table S1. Simpler models outperform geometric GNNs.
4.3 Discussion
Figure 4 and Table S1 present our benchmarking results and Figure S1 presents the corresponding parity plots. The MLP and the RF were trained on the full training set, but tested on both the full test set (in orange) and the subset of the test set that has CIF files (in red). The goal is to be able to compare their performance directly with geometric models, given that the variance of the CIF subset is larger.
The two simple models, RF and MLP, outperform all 3D geometric models both in the cross-validation and the test performance even when comparing with the subset of the test set that has CIF files. There are two factors that could explain this result. First, the RF and the MLP used the full training set of 478 structures while the other models were limited to the subset of 254 entries that have CIF files. Second, the geometric models use crystal information to infer properties of the crystal, but they do not properly handle partial occupancies which, as discussed before, are very common in SSE materials and are present in about 3/4 of our CIF files. In order to use these models without modification on our dataset we rounded occupancies to the nearest integers which can lead to important changes in the composition.
To partly verify the above claim that dataset size and the presence of partial occupancy can explain the increased performance of the simple models, we retrained them on the subset of entries that have CIF files only. Doing so, the MAE of the MLP increased to 3.15 while that of the RF was maintained at 1.87. Therefore, dataset size does seem to have a significant impact on the MLP and may explain the difference in performance between that model and the larger models. Random forest still performs well even given less data. Rounding compositions to the nearest integer on the other hand, had little effect on both the RF and the MLP. Rounding compositions is similar to rounding site occupancy, but it does not have exactly the same effect. Nevertheless, it indicates that the absence of partial occupancy likely does not explain the difference in performance between the simple models and the more complex ones.
To further explore the effects of partial occupancy, which, as explained in Section 3.1, is an important concept in this field, we introduce new implementations of both CGCNN and SO3Net (dis-CGCNN and dis-SO3Net) that take into account partial occupation (disorder). In both cases, the atomic embedding is replaced with a site embedding that is an average over the element embeddings weighted by occupancy. We trained these models using the same optimal hyperparameters as their original version. The results presented in Figure 4 and at the bottom of Table S1 show a small improvement in cross-validation performance but it does not translate into significantly better test performance.
The 3D geometric models not only performed poorly compared to simple ML models using less structural information, but their performance on the test set was barely better or sometimes worse than predicting the median of the training set (doted line in Figure 4). This shows that these large models can easily overfit small experimental datasets which was also observed in other studies [8]. Moreover, given that the cross-validation splits were chosen randomly within the training set and that the test set was build using the method described in Section 3.3, the relatively large difference in performance between the validation and testing sets illustrate the importance of carefully building leakage-free test sets and that choosing the test set randomly would have most likely led to a false impression of performance.
It is important to note that there may exist more recent GNN architectures that perform better on this task, however given the fact that some of the more recent models tested here still perform equally or close to state-of-the-art models on scalar predictive tasks [18] we do not believe these newer models would perform significantly better on OBELiX and would likely suffer from the same limitations.
[…]
It is important to bear in mind that the variability of the prediction accuracy is high in this small data regime as illustrated by the validation MAEs’ standard deviations and that much of the difference between models falls within that variability. Performance is dependent on the (random) choice of cross-validation splits which ultimately dictate the choice of hyperparameters. Complex GNNs with more hyperparameters are more prone to overfitting hyperparameters to a specific set of splits which makes them particularly difficult to tune and compare. Indeed, the variability across folds is smaller for the RF and MLP than for the GNNs.
5 Limitations
We have built OBELiX as carefully as possible making sure that all features match the measured ionic conductivity correctly. However, since data is reported and measured in very different ways across journals and decades, there most probably remains inconsistencies between some of the entries especially in terms of atomic positions which are particularly difficult to measure and report. We will continue to improve the dataset as these issues come to light.
Varying factors outside the composition and crystal structure including the measurement conditions (frequency, pressure, measuring device, metalization, etc.) and the microstructure (grain size, porosity, phase purity, etc.) that depend on the fabrication process (sintering, cold/hot press, pulverization, heat treatment, etc.) may have important effects on the measured ionic conductivity. The absence of these factors in OBELiX sets a bound to the performance of the models presented here that is partially, but not fully captured by the experimental uncertainty discussed in section 4. The repeated materials presented in Figure 3 could serve as a useful starting point to identify which of these numerous factors have the most impact on the measured conductivity and dictate what additional features could be added to the dataset.
OBELiX is small for ML standards. The difficulty of building an experimental dataset is that there is only a limited number of experiments that were actually performed. Section 4 shows how challenging it is to train existing models on such a small data regime. Ultimately, it highlights the need for models, training architectures and benchmarks tailored for small data regimes, that could benefit numerous applied fields with similarly limited experimental data (e.g. [1]). Moreover, OBELiX can be used as a tool to validate and improve molecular dynamics (MD) based methods which are widely applicable across materials science and could later serve as a way to generate a significantly larger computational database of ionic conductivity. For example, in subsequent work, we are currently using a susbet of OBELiX to compare the performance of MLFFs and ab-inito methods when predicting ionic conductivity with various MD simulation conditions. Our dataset provides an opportunity to quantitatively test the performance of MLFFs on long timescale MD simulations or ML methods such as LiFlow [21] aimed at accelerating them.
We benchmarked ionic conductivity prediction on our dataset with popular existing models as is and using standard training and hyperparameter tuning. We are aware that performance could be improved by modifying the model architectures, training procedure or with data augmentation, but we consider that these methods would not be “baselines” and are outside the scope of this paper.
6 Conclusion and outlook
In this paper, we presented OBELiX, a dataset of 599 materials with experimental room temperature ionic conductivities curated by domain experts, including 321 structures with full crystallographic information. We gathered these materials from existing databases and manually extracted data from the literature to build a consistent, easy-to-access database of solid-state electrolyte materials. We benchmarked several ML models and found that the simple random forest model had the best predictive performance. Modern geometric GNNs on the other hand, likely over-fit and were unable to perform well on our carefully designed test set. These findings highlight the immense opportunity for improvement in ML methods specific to this task and tailored for low data regimes.
We hope that OBELiX will serve as a reference point to train and test ionic conductivity models for the ML and computational materials science community in general, ultimately advancing solid-state battery technology.
7 Data availability
All data is freely available on our public repository222github.com/NRC-Mila/OBELiX as a single csv or xlsx file accompanied by a set of 321 CIF files, including 291 with added random noise. The same data is also available on Kaggle333www.kaggle.com/datasets/flixtherrien/obelix. All experiments were performed with OBELiX version 1.0.0 [30].
Supplemental Information
Appendix A Parity plots and data from Figure 4
Table S1 contains the information from Figure 4 in a tabulated form. Figure S1 presents parity plots for benchmarking experiments discussed in Section 4.
Appendix B Baseline models
We trained and tested 7 ML models which are briefly described below:
  1. 1.
  2. 2.
  3. 3.
    PaiNN [25] enforces E(3)-equivariance, enabling accurate modeling of atomic interactions and force predictions.
  4. 4.
    SchNet [27] learns continuous filter representations, making it effective for capturing atomic environments.
  5. 5.
    M3GNet [7]integrates message passing with three-body interactions, improving property predictions for crystalline materials.
  6. 6.
    SO3Net [26] leverages spherical harmonics to enhance equivariant representations for molecular and solid-state systems.
  7. 7.
    CGCNN [32] models crystal structures directly as graphs, making it a strong baseline for learning structure-property relationships.
Table S2 shows the best hyperparameter sets for each model presented in Table S1
Table S1: Benchmarking of various ML models with and without pretraining. For the median prediction, the random forest (RF) and the multilayer perceptron (MLP), results are presented for the full dataset and numbers in parenthesis are results for the subset of the test set that has CIF files. All other results apply only to entries with CIF files. "p-" indicates a model that was pretrained and "dis-" indicates a model that was modified to take partial occupancy (disorder) into account.
@@ Appendix B Baseline models
dis-CGCNN 1.52 $\pm$ 0.29 2.71
dis-SO3Net 1.72 $\pm$ 0.07 2.86
Figure S1: Parity plots for benchmarked models. a) Random Forest b) Multilayer perceptron c) PaiNN d) SchNet e) M3GNet f) SO3Net g) CGCNN h) PaiNN with pretraining i) SchNet with pretraining j) M3GNet with pretraining k) CGCNN with pretraining l) CGCNN with disorder (partial occupancy) m) SO3Net with disorder (partial occupency)
Table S2: The selected hyperparameters for the baseline models.
Model Hyperparameter Value
RF max_depth 36
max_features sqrt
min_samples_leaf 1
n_estimators 50
MLP activation relu
batch_size 16
early_stopping True
hidden_layer_sizes
[64, 64, 64,
64]
learning_rate adaptive
learning_rate_init 0.01
max_iter 1000
n_iter_no_change 100
solver adam
@@
target_property "graph"
readout_type "set2set"
@@
Appendix C Effects of added random noise
Table S3 presents the cross-validation and test MAEs for the models trained on randomized atomic positions. The models were trained on the randomized data and tested on the original CIFs. There is no significant difference in performance between models trained on the original data and models trained on data with added random noise on atomic positions.
Table S3: Performance of the 5 geometric models on the public dataset with added random noise to the atomic positions. Test results are on the original test set.
Model Cross-validation MAE Test MAE
PaiNN 2.03 $\pm$ 0.27 2.95
SchNet 1.99 $\pm$ 0.22 2.78
M3GNet 1.83 $\pm$ 0.29 2.91
SO3Net 1.98 $\pm$ 0.23 2.79
CGCNN 1.94 $\pm$ 0.42 2.95
Appendix D Computational resources used for benchmarking
Table S4 presents the resources used to find optimal hyperparameters and train each model.
Table S4: Resource usage for benchmarking
Model Hardware Hyperparameter tuning Final training
RF AMD EPYC 7502 (1 core) 7min 1s
MLP AMD EPYC 7502 (1 core) 50min 14s
PaiNN NVidia A100 GPU 2h40min 2min
SchNet NVidia A100 GPU 1h55min 2min
M3GNet NVidia A100 GPU 3h35min 2min
SO3Net NVidia A100 GPU 2h25min 2min
CGCNN NVidia A100 GPU 1h40min 1min